OpenStudy (anonymous):

I need the explanations of how as well..thanks Given the polynomial f(x) = x^4 - 37x^2+36?a. Use Descartes Rule of Signs to determine the number of positive and negative roots.b. Use the Rational Zero Theorem (aka Rational Roots Theorem) to determine a list of possible zeros.c. Use the Intermediate Value Theorem to prove that the polynomial has a zero in the interval [0, 4].d. Solve for the zeros of f(x).

OpenStudy (anonymous):

a good explanation of descartes rule of sign, as well as a couple worked out examples, is here http://www.purplemath.com/modules/drofsign.htm

OpenStudy (anonymous):

we can work this one out in detail if you like

OpenStudy (anonymous):

then, we can work out the next part Use the Rational Zero Theorem (aka Rational Roots Theorem) to determine a list of possible zeros which is absolutely stated incorrectly

OpenStudy (anonymous):

Yes please show me how to work this out in detail because I have no clue as yo where to start

OpenStudy (anonymous):

ok for descares rule of sign, lets list the coefficients of your polynomial \[f(x) = x^4 - 37x^2+36\] they are 1, -37, 36

OpenStudy (anonymous):

there are 2 changes in sign of the coefficients, from \(+1\) to \(-37\) and then from \(-37\) to \(+36\)

OpenStudy (anonymous):

is that part clear? that is, it is clear how to count the changes in sign?

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

ok so this tell you there are at most 2 POSITIVE zeros it does not say there are 2, it says there are at most 2 and you count down by 2's so there could be 2 or there could be none

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now we compute \(f(-x)=(-x)^4-37(-x)^2+36=x^4-37x^2+36=f(x)\) because this is an even function (all exponents are even) generally this is not the case, this is just a special example in this case we see there are also at most two NEGATIVE zeros so there are either 2 negative zeros , or no negative zeros

OpenStudy (anonymous):

again this is a special example because this is an even function it is symmetric with respect to the y axis, so if there are two positive zeros there are also two negative zeros. this is not always the case, it is particular to this example

OpenStudy (anonymous):

that takes care of descartes rule of sign we note the changes in sign of the coefficients of \(f(x)\) to count the number of possible positive zeros, and the changes is sign of the coefficients of \(f(-x)\) to find the number of possible negative zeros

OpenStudy (anonymous):

you with me so far?

OpenStudy (anonymous):

if so we can move on to the rational root theorem

OpenStudy (anonymous):

Yes I am

OpenStudy (anonymous):

ok before we continue, this problem has been cooked up to be very easy, because it is a quadratic in \(x^2\)and you can factor it as \((x^2-1)(x^2-36)\) so the zeros are easy to find, but lets put that on hold for a second

OpenStudy (anonymous):

Alright

OpenStudy (anonymous):

the rational root theorem tells you how to find the POSSIBLE RATIONAL roots first off no one says the roots have to be rational but IF they are rational (i.e. fractions) then they are of the form \(\frac{p}{q}\) and the rational root theorem tells you that the numerator \(p\) divides the constant and the denominator \(q\) divides the leading coefficient

OpenStudy (anonymous):

in your case the constant is 36 and the leading coefficient is 1, so your "possible rational roots" must be in this case integers that divide 36 there are a lot of them

OpenStudy (anonymous):

\[1,2,3,6,-,12,18,36\] all divide 36, as do their opposites \[-1,-2,-3,-6,-9,-12,-18,-36\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

it would be even worse if the leading coefficient was not one, because then you would have a large list of fractions as well

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

and don't forget, no one says the roots have to be rational they could be none of these

OpenStudy (anonymous):

now for part 3 check that \(f(4)\) and \(f(0)\) have opposite sign in this case \(f(0)=36\) i will let you compute \(f(4)\) for your self, but it will probably be negative therefore, since the function is positive when \(x=0\) and negative when \(x=4\) and continuous (since it is a polynomial) it must have a zero (i.e. cross the \(x\) axis, go from being positive to negative) somewhere in the interval \((0,4)\)

OpenStudy (anonymous):

now as i wrote above, finding the zeros in this case is very simple it is a polynomial of degree 4, which should make it very hard, but there is no third degree term and no first degree term, so you can think of it as a quadratic in \(x^2\) and easily factor as \[(x^2-1)(x^2-36)\] set each factor equal to zero and solve for \(x\)

OpenStudy (anonymous):

hope all steps are clear

OpenStudy (anonymous):

Thank you